OpenStudy (anonymous):

a chain is held on a frictionless table with 1/3 of its length hanging over the edge. If the chain has length L = 25 cm and mass m = 34 g, how much work is required to pull the hanging part back onto the table?

OpenStudy (anonymous):

\[\huge \int\limits_{0}^{\frac{.25 }{ 3}}\frac{ .034 }{ .25}*g*y dy\]

OpenStudy (anonymous):

the answer suppose to be in Newtons

OpenStudy (anonymous):

it is.

OpenStudy (anonymous):

see the meters and kg in that equation?

OpenStudy (vincent-lyon.fr):

No, if the answer is the work needed, the answer must be in joules, not in newtons.

OpenStudy (anonymous):

I think he meant newton meters. but whatever.